# Llama-VITS: Enhancing TTS Synthesis with Semantic Awareness

### 10 Apr 2024 v1

@@ 1.   Introduction
  • We introduce a new evaluation framework that includes traditional objective metrics for acoustic performance of synthesized speech and incorporates ESMOS, a subjective metric to assess emotional expression.
@@ 2.   Related Work
TTS technology has significantly advanced in learning acoustic features through the above structural evolution. However, comprehending and conveying semantics and emotions remain challenging. Since BERT-like LMs have demonstrated profound capabilities in understanding semantics through extensive pre-training on vast text corpora, some studies have integrated BERT-like LMs with TTS technology to enhance the expression in synthesized speech. Nonetheless, research on incorporating GPT-like LMs within TTS technology is notably scarce.
@@ 2.1.   Text-To-Speech Models
TTS task aims to generate natural, fluent, and easily comprehensible speech. Traditional TTS systems, e.g., a Statistical Parametric Speech Synthesis (SPSS) system (Taylor, 2009), usually comprise multiple distinct components. These include a frontend module that converts text into linguistic features (such as duration and pitch), an acoustic model that maps these linguistic features to acoustic features, and a vocoder responsible for generating speech waveforms from the acoustic features. Over the past decades, the complexity of traditional models has been notable, attributed to their reliance on manually engineered features and the intricate communication required between modules.
@@
The advent of end-to-end TTS models marks a significant milestone in TTS technology, increasingly reducing the distinction between synthesized speech and human voice. End-to-end models are capable of transforming raw text directly into final speech output, which not only streamlines the structural complexity of TTS systems and facilitates easier deployment but also significantly reduces the dependency on manual feature engineering, simplifying the training process. Moreover, they notably enhance the naturalness and intelligibility of the speech, thereby becoming the predominant architecture in TTS models. For instance, Char2Wav (Sotelo et al., 2017) introduces an attentive encoder-decoder framework for direct speech synthesis from text input. Tacotron (Wang et al., 2017) undertakes training from the ground up and directly predicts linear spectrograms. Furthermore, the speech produced by Tacotron2 (Shen et al., 2018) closely mirrors the natural human voice.
@@
In the realm of end-to-end TTS models, many have adopted a non-autoregressive model architecture. This architecture enables parallel data processing, where the model’s output generation does not depend on the output of the previous time step, thereby enhancing processing speed. It also circumvents the error accumulation issue inherent in traditional autoregressive models during audio generation, which significantly boosts TTS performance. FastSpeech (Ren et al., 2019) and its variants exemplify this trend. FastSpeech employs a transformer-based architecture to generate mel-spectrograms in parallel. Building on FastSpeech, FastPitch (Łańcucki, 2021) predicts pitch contours during inference, enabling the production of more expressive and high-quality speech. FastSpeech2 (Ren et al., 2022) further refines the model by incorporating explicit duration prediction and introducing pitch and energy as conditional inputs for optimization.
@@
Previous non-autoregressive approaches typically involve distinct training phases for acoustic models and vocoders. However, VITS (Kim et al., 2021) introduces a more natural-sounding output compared to these two-stage systems through its one-stage parallel end-to-end architecture. Innovatively, VITS incorporates variational inference combined with normalizing flows and employs an adversarial training methodology. Due to VITS’s exemplary performance across multiple TTS benchmarks, it has been selected as the foundational TTS model for our system.
2.2.   Fine-tuning BERT-like models for TTS
@@
While TTS models have increasingly advanced in replicating acoustic features, insufficient training data can hinder the model’s ability to learn the emotional nuances of the same input across different contexts, thus limiting its expressive capabilities. Consequently, researchers have turned to leveraging the robust transfer learning capabilities of BERT-like models. Ultimately, TTS systems that incorporate pre-trained and fine-tuned BERT-like models have achieved better understandings of semantics and enhanced generated speech, marking a significant advancement in TTS technology.
@@
Hayashi et al. (2019) utilized a pre-trained BERT model as an auxiliary input to enhance a Tacotron2-based TTS system, resulting in improved speech naturalness. Similarly, Yang et al. (2019) applied a pre-trained BERT to achieve enhanced front-end accuracy. Kenter et al. (2020) demonstrated that integrating a BERT model, pre-trained on extensive unlabeled data and fine-tuned for speech, into an RNN-based speech synthesis system enhances prosody. Kenter et al. (2020) specifically suggest updating the BERT network’s parameters during the training of their RNN-based speech synthesis model, emphasizing the critical role of fine-tuning the BERT component for optimal outcomes. As prompt tuning draws wide attention in guiding text or image generation, PromptTTS (Guo et al., 2022) takes a prompt with both style and content descriptions as input to synthesize the corresponding speech. Specifically, PromptTTS consists of a style encoder and a content encoder to extract the corresponding representations from the prompt, and a speech decoder to synthesize speech according to the extracted style and content representations. Experiments show that PromptTTS can generate speech with precise style control and high speech quality.
@@ 2.3.   Integrating GPT-like models for TTS
However, research on leveraging GPT-like models to aid semantic understanding in TTS systems is very limited. For instance, Stephenson et al. (2021) explores the potential of improving speech synthesis prosody with the GPT model by modifying text input. Such an approach potentially restricts TTS applications, as altering the input is often undesirable. Furthermore, the findings were not particularly significant.
@@
Related research that employs Llama2 in speech and other multimodal tasks (Radhakrishnan et al., 2023; Zhang et al., 2023), coupled with the ongoing efforts to reduce computing costs associated with Llama2222https://huggingface.co/4bit/Llama-2-70b-chat-hf, underscores the model’s significant research interest and its promising prospects in multimodal applications.
@@ 3.1.   Semantic Embeddings Derived from Llama2
In the [EIS_Word] and [EIS_Sentence] strategies, unlike the above approaches that utilize the sentence itself for representation, we derive the semantic representation of sentence $s$ based on Llama2’s comprehension $u$. Specifically, we employ prompts, as illustrated in 1(a) and 1(b), respectively, to obtain Llama2’s understanding of sentence $s$ in terms of Emotion, Intention, and Speaking Style, denoted as $u$, and calculate the average of this understanding’s representation to serve as the semantic token.
@@ 4.1.   Experimental Settings
We utilized full LJSpeech, 1-hour LJSpeech, and EmoV_DB_bea_sem dataset for our experimental verification. LJSpeech  [^3] https://keithito.com/LJ-Speech-Dataset/ comprises 24 hours recorded of English speech by single female speaker, where we evaluate how the tokens extracted from Llama2 can help improve the speech naturalness. Besides the full LJSpeech dataset, we also randomly filtered 1-hour LJSpeech which contains only 1-hour records as an ablation study to show how dataset size influences the results. EmoV_DB (Adigwe et al., 2018) is a database of emotional speech that contains data for male and female actors in English and a male actor in French. EmoV_DB covers 5 emotion classes, amused, angry, disgusted, neutral, and sleepy. To factor out the effect of different speakers, we filtered the original EmoV_DB dataset into the speech of a specific female English speaker, bea. Then we use Llama2 to predict the emotion label of the transcript chosen from the above 5 emotion classes, and select the audio samples which has the same predicted emotion. The filtered dataset contains 22.8-min records for training. We named the filtered dataset EmoV_DB_bea_sem and investigated how the semantic tokens from Llama2 behave in naturalness and expressiveness on this filtered dataset. Refer to Appendix Appendix A. Dataset Statistics for more dataset statistics.
@@
In evaluating ESMOS, we randomly chose 5 samples from the total 51 test samples proportionally divided by us and received 100 test results from different speakers on Amazon Mechanical Turk. The result significance level is thus 500. Each participant is asked to give a score on emotion similarity compared with ground truth in a 5-scale: Excellent Match 5, Good Match 4, Fair Match 3, Poor Match 2, Bad Match 1888Note that in the ESMOS experiments, participants are asked to ignore the speakers’ voice, style, and audio quality and only consider the emotiveness of the speech..
@@ 6.   Discussions
In this section, we discuss factors influencing current outcomes. Based on this discussion, we also point out the directions for future work in Appendix Appendix B. Future Work.
@@ 11.   Language Resource References
Appendix A. Dataset Statistics
@@
Appendix B. Future Work

### 12 Apr 2024 v2

@@ 1.   Introduction
  • We quantitatively verified our findings using both subjective and objective metrics.
@@ 2.   Related Work
TTS technology has significantly advanced in learning acoustic features through structural evolution. However, comprehending and conveying semantics remain challenging. Since BERT-like LMs have demonstrated profound capabilities in understanding semantics through extensive pre-training on vast text corpora, some studies have integrated BERT-like LMs with TTS technology to enhance synthesized speech. Nonetheless, research on incorporating GPT-like LMs within TTS technology is notably scarce.
@@ 2.1.   Text-To-Speech Models
TTS task aims to generate natural, fluent, and easily comprehensible speech. Traditional TTS systems, e.g., a Statistical Parametric Speech Synthesis (SPSS) system (Taylor, 2009), usually comprise multiple distinct components. These include a frontend module that converts text into linguistic features (such as duration and pitch), an acoustic model that maps these linguistic features to acoustic features, and a vocoder responsible for generating speech waveforms from the acoustic features. Over the past decades, the complexity of traditional models has been notable, attributed to their reliance on manually engineered features and the intricate communication between modules.
@@
The advent of end-to-end TTS models marks a significant milestone, increasingly reducing the distinction between synthesized speech and human voice. End-to-end models are capable of transforming raw text directly into final speech output, which not only streamlines the structural complexity of TTS systems and facilitates easier deployment but also significantly reduces the dependency on manual feature engineering, simplifying the training process. Moreover, they notably enhance the naturalness and intelligibility of the speech, thereby becoming the predominant architecture in TTS models. For instance, Char2Wav (Sotelo et al., 2017) introduces an attentive encoder-decoder framework for direct speech synthesis from text input. Tacotron (Wang et al., 2017) undertakes training from the ground up and directly predicts linear spectrograms. Furthermore, the speech produced by Tacotron2 (Shen et al., 2018) closely mirrors the natural human voice.
@@
In the realm of end-to-end TTS models, many have adopted a non-autoregressive architecture. This architecture enables parallel data processing, where the model’s output generation does not depend on the output of the previous time step, thereby enhancing processing speed. It also circumvents the error accumulation issue inherent in traditional autoregressive models, which significantly boosts TTS performance. FastSpeech (Ren et al., 2019) and its variants exemplify this trend. FastSpeech employs a transformer-based architecture to generate mel-spectrograms in parallel. Building on FastSpeech, FastPitch (Łańcucki, 2021) predicts pitch contours during inference, enabling the production of more expressive and high-quality speech. FastSpeech2 (Ren et al., 2022) further incorporates explicit duration prediction and introduces pitch and energy as conditional inputs.
@@
Previous non-autoregressive approaches typically involve distinct training phases for acoustic models and vocoders. VITS (Kim et al., 2021) introduces a more natural-sounding output compared to these two-stage systems through its one-stage parallel end-to-end architecture. Innovatively, VITS incorporates variational inference combined with normalizing flows and employs an adversarial training methodology. Due to VITS’s exemplary performance across multiple benchmarks, we select it as the foundational TTS model for our system.
2.2.   Fine-tuning BERT-like LMs for TTS
@@
While TTS models have increasingly advanced in replicating acoustic features, insufficient training data can hinder the model’s ability to learn the semantic nuances of the same input across different contexts, thus limiting its expressiveness. Consequently, researchers have turned to leveraging the transfer learning capabilities of BERT-like LMs. Ultimately, TTS systems that incorporate pre-trained and fine-tuned BERT-like LMs have achieved better understandings of semantics and enhanced generated speech, marking a significant advancement.
@@
Hayashi et al. (2019) utilized a pre-trained BERT model as an auxiliary input to enhance a Tacotron2-based TTS system, resulting in improved speech naturalness. Similarly, Yang et al. (2019) applied a pre-trained BERT model to achieve enhanced front-end accuracy. Kenter et al. (2020) demonstrated that integrating a BERT model, pre-trained on extensive unlabeled data and fine-tuned for speech, into an RNN-based TTS system enhances prosody. Kenter et al. (2020) specifically suggest updating the BERT’s parameters during the training of their RNN-based speech synthesis model, emphasizing the critical role of fine-tuning the BERT component for optimal outcomes. As prompt tuning draws wide attention in guiding text or image generation, PromptTTS (Guo et al., 2022) takes a prompt representation with both style and content descriptions from a BERT model as input to generate speech with precise style control and high speech quality.
@@ 2.3.   Integrating GPT-like LMs for TTS
However, research on leveraging GPT-like models to aid TTS systems is very limited. Stephenson et al. (2021) explores the potential of improving speech synthesis naturalness by text input lookahead with GPT prediction. Such an approach potentially restricts TTS applications, as altering the input is often undesirable. Furthermore, the findings were not verified by human subjective evaluation. Saito et al. (2023) suggest employing ChatGPT [^1] https://openai.com/blog/chatgpt to aid in empathetic dialogue speech synthesis by extracting the context of conversations. They particularly instruct ChatGPT to produce three keywords that encapsulate the intention, emotion, and speaking Style of speech observed in the dialogue history. These keywords are subsequently utilized to train a speech synthesis model. However, due to the inaccessibility of ChatGPT to the public, the researchers resort to processing ChatGPT’s outputs with BERT to extract embeddings. This approach essentially positions ChatGPT as an alternative to manual annotation, yet it does not delve into investigating ChatGPT’s internal representations and their potential impact on speech-related tasks.
@@
Related research that employs Llama2 in speech and other multimodal tasks (Radhakrishnan et al., 2023; Zhang et al., 2023), coupled with the ongoing efforts to reduce computing costs associated with Llama2333https://huggingface.co/4bit/Llama-2-70b-chat-hf, underscores the model’s significant research interest and its promising prospects in multimodal applications.
@@ 3.1.   Semantic Embeddings Derived from Llama2
In the [EIS_Word] and [EIS_Sentence] strategies, unlike the above approaches that utilize the sentence itself for representation, we derive the semantic representation of sentence $s$ based on Llama2’s comprehension $u$. Adapted from Saito et al. (2023)’s practice, we employ prompts as illustrated in 1(a) and 1(b), respectively, to obtain Llama2’s understanding of sentence $s$ in terms of Emotion, Intention, and speaking Style, denoted as $u$, and calculate the average of this understanding’s representation to serve as the semantic embedding.
@@ 4.1.   Experimental Settings
We utilized full LJSpeech, 1-hour LJSpeech, and EmoV_DB_bea_sem dataset for experimental verification. LJSpeech [^4] https://keithito.com/LJ-Speech-Dataset/ comprises 24 hours recorded of English speech by single female speaker, where we evaluate how the embeddings extracted from Llama2 can help improve the speech naturalness. Besides full LJSpeech dataset, we also randomly filtered 1-hour LJSpeech which contains only 1-hour records as an ablation study to show how dataset size influences. EmoV_DB [^5] https://www.openslr.org/115/ (Adigwe et al., 2018) is a database of emotional speech that contains data for male and female actors in English and French. EmoV_DB covers 5 emotion classes, amused, angry, disgusted, neutral, and sleepy. To factor out the effect of different speakers, we filtered the original EmoV_DB dataset into the speech of a specific female English speaker, bea. Then we use Llama2 to predict the emotion label of the transcript chosen from the above 5 emotion classes, and select the audio samples which has the same predicted emotion. The filtered dataset contains 22.8-min records for training. We named the filtered dataset EmoV_DB_bea_sem and investigated how the semantic embeddings from Llama2 behave in naturalness and expressiveness on it. Please refer to Appendix A 12 for more dataset statistics.
@@
In evaluating ESMOS, we randomly chose 5 samples from the total 51 test samples proportionally divided by us and received 100 test results from different speakers on Amazon Mechanical Turk. The result significance level is thus 500. Each participant is asked to give a score on emotion similarity compared with ground truth in a 5-scale: Excellent Match 5, Good Match 4, Fair Match 3, Poor Match 2, Bad Match 1101010Note that in the ESMOS experiments, participants are asked to ignore the speakers’ voice, style, and audio quality and only consider the emotiveness of the speech..
@@ 6.   Discussions
In this section, we discuss factors influencing current outcomes. Based on this discussion, we also point out the directions for future work in Appendix 13.
@@ 10.   Bibliographical References
  • Saito et al. (2023) Yuki Saito, Shinnosuke Takamichi, Eiji Iimori, Kentaro Tachibana, and Hiroshi Saruwatari. 2023. Chatgpt-edss: Empathetic dialogue speech synthesis trained from chatgpt-derived context word embeddings.
@@ 11.   Language Resource References
12.   Appendix A. Dataset Statistics
We have summarized the dataset statistics in the table below.
@@
13.   Appendix B. Future Work